# Deep Reinforcement Learning for Dynamic Order Picking in Warehouse Operations

### v1

@@ 5.1 Benchmark Algorithms
To evaluate our order-picker routing algorithm, we utilize two well-known benchmark studies from the literature:
@@
    Lu et al. (2016): This study adapts the static order picking approach for dynamic environments, under certain assumptions. Specifically, the order picker begins a tour when incoming orders reach 25% of device capacity and can be re-routed if new orders arrive within capacity limits. Furthermore, assuming narrow cross-aisles, the picker cannot be re-routed while traveling within them. This method is referred to as the Interventionist Routing Algorithm (IRA). The authors have demonstrated the superiority of their method compared to an effective heuristic known as the largest gap heuristic.
We note that the source code for the aforementioned algorithms is unavailable. More importantly, the core of the above-mentioned studies relies on Dynamic Programming (DP) methods, which are favored for their efficacy in picker routing. However, these DP methods operate under certain assumptions, and extending them to accommodate the diverse routing scenarios found in real-world warehouse operations is not a trivial task. For instance, although picker re-routing within cross-aisles is practical and significantly impacts operational efficiency (as will be shown in the numerical results), it is assumed to be prohibited in these DP methods. Note that our proposed DRL agent automatically learns to apply cross-aisle re-routing as needed. Therefore, to replicate the results of the aforementioned studies under their assumptions, as well as to make these methods adaptable for various routing scenarios, we replace DP with a generic TSP solver in their methods to compute optimal routes. This allows for more flexible routing scenarios, such as using smaller initial pick list sizes or permitting cross-aisle re-routing. Although a generic TSP solver is slower, it does not alter the fundamental nature of the algorithms’ performance. We exclude the computational time of the generic TSP solver in our experiments. Additionally, in our computational study, we do not report any computational time, as the proposed DRL method makes decisions instantaneously at any given state. Thus, the runtime is negligible during testing, with all computational time spent during the training phase. With these considerations, we establish five baseline models for our computational study:
@@
These variants serve as additional benchmarks to assess the effects of initial pick list size and cross-aisle re-routing constraints on overall picking performance. Note that all baseline comparisons utilize a first-come-first-served batching policy, consistent with the above referenced benchmark studies. A comparison of different baseline assumptions is listed in Table 1.
@@ 5.2 Training and Testing Details
The proposed DRL approach was implemented in Python 3.9 using PyTorch 1.11. The hyperparameter configuration consisted of a batch size of 64, with training conducted over 4,500 episodes, each comprising 1,000 steps. The neural network architecture consisted of initial layers with dimensions $H^{o}=160$ and $H^{p}=64$, followed by the fully connected layers with dimensions $F_{1}=256,F_{2}=128$ and $F_{3}=64$. Experience replay was employed with a memory capacity of $C=200,000$. The optimization process utilized a learning rate of 0.0001, discount factor $\gamma=0.99$, and target update coefficient $\tau=0.001$. The reward function, $R$, is defined as the sum of the number of storage locations in each aisle and the total number of aisles, ensuring a positive reward for successful item pickup regardless of the agent’s initial location.
We note that a DRL model trained for a specific order arrival rate may not generalize well to different arrival rates. Machine learning frameworks often exhibit sensitivity to the training distribution and its parameter values (Tsotsos et al., 2019; Sobhanan et al., 2023). So, we conduct a series of training experiments with varying arrival rates: $\lambda\in\{0.02,0.04,0.06,0.08\}$. Furthermore, we investigate the trade-off between prioritizing distance minimization and order completion time by adjusting the weight hyperparameter $\alpha\in\{0.1,0.5,1.0\}$. This results in the training of 12 models, given that there are four values for $\lambda$ and three values for $\alpha$. The training process for each model takes approximately 7 hours to complete on a MacBook Pro equipped with an Apple M1 Chip and 16GB RAM. The training curve for one of the models, with $\lambda=0.08$ and $\alpha=1$, is shown in Figure 3 as an example. The other models follow a similar convergence pattern.
Figure 3: Training curve showing convergence of 20-period moving average reward for $\lambda=0.08$ and $\alpha=1$
@@
We evaluate the performance of each trained model using test instances with varying values of $\lambda$ to assess their generalizability across different arrival patterns and optimization objectives. For this testing, we use 9 distinct values for $\lambda$, ranging from $0.01$ to $0.09$ in increments of $0.01$, and evaluate each of the 12 trained models against all of these values. Each model is tested against 1 in-sample scenario and 8 out-of-sample scenarios. Note that test scenarios with $\lambda\in\{0.01,0.03,0.05,0.07,0.09\}$ serve as out-of-sample scenarios for all trained models. As previously mentioned, each testing scenario consists of 10 simulation runs, with each run representing an 8-hour work shift.
@@
Figure 4 illustrates a pick cycle recommended by our trained DRL agent for a specific scenario with $\lambda=0.06$ and $\alpha=1$. In this context, the term ‘time’ refers to the elapsed time since the start of the work shift (not the computational time). Accordingly, the picker begins this cycle 661 seconds into the work shift with a full capacity of $K=20$. The blue square represents the depot, while the red square denotes the picker with its remaining capacity highlighted. Since the horizontal distance within aisles is negligible, orders within storage locations on both sides are shown as green dots inside the aisles. At 673 seconds, as depicted in Figure 4(a), the order picker arrives at aisle 9 and retrieves an item from the second row. Simultaneously, a new demand arises for an item on the third row in the same aisle. The picker efficiently fulfills all three pick requests in aisle 9 before moving sequentially to aisles 7 and 5, as indicated by the red line. Figure 4(b) illustrates the picker carrying five items with a remaining capacity of 15 units as it proceeds towards aisle 6 to pick up the next set of requested items. At this point, two new orders appear at position 13 in aisle 6. The DRL agent’s demand forecasting guided the picker to this aisle. Figure 4(c) shows the picker completing pick-ups in aisle 6 before returning to the depot via the same path. In Figure 4(d), the picker concludes the drop-offs at the depot, where each drop-off takes one second. This marks the end of the pick cycle, and the picker is prepared for the next assignment. Note that new orders have arrived at the warehouse at various time steps during the picker’s traversal. Throughout the pick cycle, the DRL agent dynamically prioritizes orders to minimize the average order fulfillment time.
@@ 5.3 Numerical Results
The results of our experiments are summarized in Tables 2, 3, and 4. Table 2 presents the average total distance traveled (ATDO) resulting from each model decisions for each instance type. The corresponding results for average order completion time (AOCT) and percentage of unfulfilled orders (PUO) are shown in Tables 3 and 4, respectively. Across these tables, the shaded cells highlight either the most desirable result or the most effective DRL model. Specifically, the trained model demonstrating the best average order completion time for each order arrival rate is highlighted. Next, we analyse the results and make a few observations.
The first observation is that, unsurprisingly, Baseline 1 results in the shortest distance traveled due to its strategy of waiting for the pick list to reach pick device capacity. This was expected, as by design, this approach allows the picker to follow the shortest possible path and move uninterrupted even when new customer orders arrive. However, as previously mentioned, the most crucial metric for optimizing warehouse operations is not the shortest distance traveled but the order completion times. Unfortunately, Baseline 1 has the worst performance in terms of average order completion times across all arrival rates and, consequently, the highest percentage of unfilled orders. Therefore, among the baselines, the variations of IRA—namely Baselines 2 through 4—are certainly better choices than Baseline 1.
@@
6 Conclusions
@@
In this research, we propose a Deep Reinforcement Learning (DRL) model designed to solve the dynamic order picking problem within a single-block warehouse layout served by an autonomous picking device. Our results demonstrate that the model’s ability to forecast customer demands and adaptively learn efficient order picking policies leads to significant enhancements in overall warehouse operations. Furthermore, the additional experiments demonstrate the robustness of our proposed model through its strong performance on out-of-sample test instances with varying order arrival rates. This work provides a foundation for future research to explore more complex warehouse environments. In particular, the extension of this DRL framework to multi-block layouts and scenarios involving multiple coordinated picking devices presents promising avenues for further advancing the field of warehouse automation. By leveraging the forecasting and adapting capabilities of DRL, warehouse operations can be both efficient and responsive to the dynamic nature of customer demands.

### v2

@@ 5.1 Benchmark Algorithms
To evaluate our order-picker routing algorithm, we employ a total of seven benchmark approaches. Two of these are widely used heuristic picking policies from the literature, specifically the S-shape and Largest-gap methods, which we discussed earlier in Section 2. These serve to analyze the solution quality of our algorithm on the problem at hand. The remaining five benchmarks are based on different configurations of approaches introduced in two well-known studies from the literature:
@@
    Lu et al. (2016): This study adapts the static order picking approach for dynamic environments, under certain assumptions. Specifically, the order picker begins a tour when incoming orders reach 25% of device capacity and can be re-routed if new orders arrive within capacity limits. Furthermore, assuming narrow cross-aisles, the picker cannot be re-routed while traveling within them. This method is referred to as the Interventionist Routing Algorithm (IRA). The authors have demonstrated the superiority of their method compared to an effective heuristic known as the largest-gap heuristic.
We note that the source code for the aforementioned algorithms is unavailable. More importantly, the core of the above-mentioned studies relies on Dynamic Programming (DP) methods, which are favored for their efficacy in picker routing. However, these DP methods operate under certain assumptions, and extending them to accommodate the diverse routing scenarios found in real-world warehouse operations is not a trivial task. For instance, although picker re-routing within cross-aisles is practical and significantly impacts operational efficiency (as will be shown in the numerical results), it is assumed to be prohibited in these DP methods. Note that our proposed DRL agent automatically learns to apply cross-aisle re-routing as needed. Therefore, to replicate the results of the aforementioned studies under their assumptions, as well as to make these methods adaptable for various routing scenarios, we replace DP with a generic TSP solver in their methods to compute optimal routes. This allows for more flexible routing scenarios, such as using smaller initial pick list sizes or permitting cross-aisle re-routing. Although a generic TSP solver is slower, it does not negatively impact the algorithm’s decision-making process. This is because we exclude the computational time of the generic TSP solver in our experiments. In other words, during the simulation of work shifts, we assume the solution time of these generic solvers is zero and do not include it in the decision-making process. Additionally, in our computational study, we do not report any computational time, as the proposed DRL method makes decisions instantaneously at any given state. Thus, the runtime is negligible during testing, with all computational time spent during the training phase. With these considerations, the remaining five baseline models for our computational study, based on optimal TSP routing, can be expressed as follows:
@@
These variants serve as additional benchmarks to assess the effects of initial pick list size and cross-aisle re-routing constraints on overall picking performance. Note that all baseline comparisons utilize a first-come-first-served batching policy, consistent with the above referenced benchmark studies. A comparison of different baseline assumptions is listed in Table 3. We note that in this table, we report only one configuration of our proposed method, as only this configuration’s results will be presented throughout this section. However, it is important to note that all the configurations used for Baselines 1–5 can also be applied to our proposed method. For interested readers, a comprehensive analysis of different variations of our proposed method is provided in B.
@@ 5.2 Training and Testing Details
The proposed DRL approach is implemented in Python 3.9 using PyTorch 1.11. Interested readers can find our codes available on https://github.com/Sasanm88/DRL-Order-Picking. The hyperparameter selection was conducted using a grid search, where the optimal value was determined based on the reward function. Table 4 presents the results of the grid search for hyperparameters. The selected configuration consists of a batch size of 64, with training conducted over 4,500 episodes, each comprising 1,000 steps. The neural network architecture consists of initial layers with dimensions $H^{o}=160$ and $H^{p}=64$, followed by the fully connected layers with dimensions $F_{1}=256,F_{2}=128$ and $F_{3}=64$. Note that $H^{o}$ and $H^{p}$ represent two components of a single layer, meaning that the structure $[H^{o},H^{p},F_{1},F_{2},F_{3}]$ effectively constitutes four layers in total. Experience replay is employed with a memory capacity of $C=200,000$. The optimization process utilizes a learning rate of 0.0001, discount factor $\gamma=0.99$, and target update coefficient $\tau=0.001$. The reward function, $R$, is defined as the sum of the number of storage locations in each aisle and the total number of aisles, ensuring a positive reward for successful item pickup regardless of the agent’s initial location.
Table 4: Selected Training Hyperparameters Using Grid Search for the DRL Model
Hyperparameter Grid Selected Value
No. of hidden layers [1,2,3,4,5] 4
Hidden layer sizes [[64, 32, 128, 256, 128], [128, 128, 128, 64, 32], [160, 64, 256, 128, 64]
$[H^{o},H^{p},F_{1},F_{2},F_{3}]$ [128, 64, 64, 256, 128], [160, 64, 256, 128, 64]]
Activation [ReLU, Tanh] ReLU
Output activation [Sigmoid, Softmax, Linear] Softmax
Optimizer [Adam, SGD] Adam
Batch size [32, 64, 128, 256] 64
We note that a DRL model trained for a specific order arrival rate may not generalize well to different arrival rates. Machine learning frameworks often exhibit sensitivity to the training distribution and its parameter values (Tsotsos et al., 2019; Sobhanan et al., 2024). So, we conduct a series of training experiments with varying arrival rates: $\lambda\in\{0.02,0.04,0.06,0.08\}$. Furthermore, we investigate the trade-off between prioritizing distance minimization and order completion time by adjusting the weight hyperparameter $\alpha\in\{0.1,0.5,1.0\}$. This results in the training of 12 models, given that there are four values for $\lambda$ and three values for $\alpha$. The training process for each model takes approximately 7 hours to complete on a MacBook Pro equipped with an Apple M1 Chip and 16GB RAM. The training curves for each trained model with varying $\lambda$ and $\alpha$ are shown in Figure 4. These figures illustrate the convergence of the training procedure. The training curves indicate that the agent tends to engage in exploration when there are no significant changes in the average reward over multiple episodes.
(a) $\lambda=0.02$
(b) $\lambda=0.04$
(c) $\lambda=0.06$
(d) $\lambda=0.08$
Figure 4: Training curves showing the convergence of the 20-period moving average reward for each trained model
@@
We evaluate the performance of each trained model using test instances with varying values of $\lambda$ to assess their generalizability across different arrival patterns and optimization objectives. For this testing, we use 9 distinct values for $\lambda$, ranging from $0.01$ to $0.09$ in increments of $0.01$, and evaluate each of the 12 trained models against all of these values. Each model is tested against 1 in-sample scenario and 8 out-of-sample scenarios. Note that test scenarios with $\lambda\in\{0.01,0.03,0.05,0.07,0.09\}$ serve as out-of-sample scenarios for all trained models. As previously mentioned, each testing scenario consists of 10 simulation runs, with each run representing an 8-hour work shift. It is worth mentioning that in addition to the experiments presented in the main body of the paper, we also conducted supplementary simulations, presented in C, where the arrival rate varies over the course of the 8-hour shift. The performance of the proposed algorithms under these time-varying arrival patterns was consistent with the results reported in the main analysis, where a constant arrival rate was assumed throughout the shift. These findings further support the robustness of the proposed models.
@@
Interested readers may refer to a YouTube animation video (https://youtu.be/LumE7YE-JMo) that illustrates the real-time progress of the DRL agent. However, in this section, we present a few representative snapshots to highlight key aspects of the agent’s behavior over time. Specifically, figure 5 illustrates a pick cycle recommended by our trained DRL agent for a specific scenario with $\lambda=0.06$ and $\alpha=1$. In this context, the term ‘time’ refers to the elapsed time since the start of the work shift (not the computational time). Accordingly, the picker begins this cycle 661 seconds into the work shift with a full capacity of $K=20$. The blue square represents the depot, while the red square denotes the picker with its remaining capacity highlighted. Since the horizontal distance within aisles is negligible, orders within storage locations on both sides are shown as green dots inside the aisles. At 673 seconds, as depicted in Figure 5(a), the order picker arrives at aisle 9 and retrieves an item from the second row. Simultaneously, a new demand arises for an item on the third row in the same aisle. The picker efficiently fulfills all three pick requests in aisle 9 before moving sequentially to aisles 7 and 5, as indicated by the red line. Figure 5(b) illustrates the picker carrying five items with a remaining capacity of 15 units as it proceeds towards aisle 6 to pick up the next set of requested items. At this point, two new orders appear at position 13 in aisle 6. The DRL agent’s demand forecasting guided the picker to this aisle. Figure 5(c) shows the picker completing pick-ups in aisle 6 before returning to the depot via the same path. In Figure 5(d), the picker concludes the drop-offs at the depot, where each drop-off takes one second. This marks the end of the pick cycle, and the picker is prepared for the next assignment. Note that new orders have arrived at the warehouse at various time steps during the picker’s traversal. Throughout the pick cycle, the DRL agent dynamically prioritizes orders to minimize the average order fulfillment time.
@@ 5.3 Numerical Results
The results of our experiments are summarized in Tables 5, 6, and 7. Table 5 presents the average total distance traveled (ATDO) resulting from each model decisions for each instance type. The corresponding results for average order completion time (AOCT) and percentage of unfulfilled orders (PUO) are shown in Tables 6 and 7, respectively. Across these tables, the shaded cells highlight either the most desirable result or the most effective DRL model. Specifically, the trained model demonstrating the best average order completion time for each order arrival rate is highlighted. Next, we analyse the results and make a few observations.
The first observation is that, unsurprisingly, Baseline 1 results in the shortest distance traveled due to its strategy of waiting for the pick list to reach pick device capacity. This was expected, as by design, this approach allows the picker to follow the shortest possible path and move uninterrupted even when new customer orders arrive. However, as previously mentioned, the most crucial metric for optimizing warehouse operations is not the shortest distance traveled but the order completion times. Unfortunately, Baseline 1 has the worst performance in terms of average order completion times across all arrival rates and, consequently, the highest percentage of unfilled orders. Therefore, among the baselines, the variations of IRA—namely Baselines 2 through 5—are certainly better choices than Baseline 1. Notably, heuristic policies such as S-shape and Largest-gap consistently perform poorly compared to their optimal distance-routing counterpart, Baseline 5, across all performance metrics. This outcome is unsurprising, as the routing decisions of these heuristics are suboptimal compared to their corresponding exact baselines, which rely on solving the TSP to optimality at all times.
@@
(a) Effect of the number of storage locations per aisle when the number of aisles is 10
(b) Effect of the number of aisles when the number of storage locations per aisle is 15
Figure 6: Training time of the DRL model for $\lambda=0.08$ and $\alpha=1$ across different warehouse configurations
6 Final Remarks
@@
In this research, we proposed a DRL model designed to solve the dynamic order picking problem within a single-block warehouse layout served by an autonomous picking device. Our results demonstrate that the model’s ability to forecast customer demands and adaptively learn efficient order picking policies leads to significant enhancements in overall warehouse operations. We benchmarked the proposed model against optimal static algorithms and widely used heuristics designed for picking distance optimization, and showed that our approach achieves a substantial reduction in average order completion times while fulfilling a significantly higher number of orders during the work shift. Furthermore, additional experiments validated the robustness of the model, as evidenced by its strong performance in out-of-sample test instances with varying order arrival rates. Our findings highlighted the adaptability and reliability of the DRL framework and established it as a promising method for advancing warehouse automation.
As mentioned, the proposed method has been shown to be suitable for single-block warehouses with a single autonomous picking device. However, even within its domain of applicability, it is important to highlight that some computational challenges may arise depending on the structure of the warehouse. Specifically, we demonstrated the effectiveness of our approach on warehouses with 10 aisles and 15 storage locations per side of each aisle, which is a common assumption in the existing body of literature. If the size of the warehouse changes, the proposed method remains effective but is expected to require significantly more computational power during the training phase. This is illustrated in Figure 6, where the computational time required to train the DRL model under varying warehouse configurations is presented with $\lambda=0.08$ and $\alpha=1$. Figure 6(a) examines the impact of increasing the number of storage locations per aisle while keeping the number of aisles fixed at 10. Conversely, Figure 6(b) explores the effect of increasing the number of aisles, with the number of storage locations per aisle side fixed at 15. In both scenarios, the results reveal a clear trend: training time increases substantially as warehouse size grows. Despite this computational cost, the DRL model demonstrates significant potential for practical problem-solving. Most warehouse configurations remain fixed once determined, making the one-time training overhead manageable. To further alleviate this computational burden, parallel computing techniques could be leveraged to distribute training across multiple GPUs or processing units.
It is also worth mentioning that this work provides a foundation for future research to explore more complex warehouse environments. While this study focuses on a single-block warehouse configuration, extending the model to multi-block layouts offers the opportunity to address the greater complexities of real-world operations. Similarly, extending the current scenario from a single autonomous picking device to one involving multiple coordinated devices, whether fully autonomous or incorporating human-robot interactions, presents a compelling avenue for further exploration. In both of these research avenues, employing parallelization and exploring suitable neural network architectures and reward functions will emerge as the primary challenges.
Appendix A Rationale for Feature-Based State Representation
A significant challenge in dynamic order-picking tasks is managing the extremely high-dimensional raw data, such as exact picker coordinates and detailed item-level positions for every storage location. To overcome this, we propose an aggregated feature-based state representation that retains all decision-relevant information while avoiding unnecessary fine-grained spatial details. Our representation captures essential routing information, picker constraints, and order demands through targeted aggregation.
We aggregate orders at the aisle level by separately considering upward and downward traversal directions within each aisle. Specifically, each aisle is represented by two aggregated indices corresponding to the potential rewards of traversing upward and downward. Potential rewards are calculated dynamically, considering both the demand (number of items) and their distance from the picker’s current position. Items closer to the picker have higher reward contributions, thereby guiding the agent to prioritize more immediately accessible orders.
The picker’s state is represented concisely through three main components:
  • Horizontal Position $(S^{H}_{t})$: Encodes the picker’s horizontal location as either in the back cross-aisle $(S^{H}_{t}=-1)$, within a pick aisle $(S^{H}_{t}=0)$, or in the front cross-aisle $(S^{H}_{t}=1)$. This abstract encoding preserves essential movement constraints without the complexity of precise coordinates.
  • Vertical Position $(S^{V_{1}}_{t},S^{V_{2}}_{t})$: Each aisle $n$ is associated with two distinct indices, $12n-$ 1 and $2n$, which align with upward and downward traversal rewards, respectively. When the picker is positioned in aisle $n$, we set $(S^{V_{1}}_{t},S^{V_{2}}_{t})=(2n-1,2n)$. This indexing scheme directly aligns picker positions with the order state representation, enabling efficient computation of aisle-specific potential rewards.
  • Remaining Capacity $(S^{C}_{t})$: Indicates the available capacity of the picker, ranging from fully loaded $(0)$ to completely empty $(K)$. Including capacity explicitly allows the agent to plan efficient routes, avoid unnecessary visits to aisles when capacity is insufficient, and effectively schedule depot returns. Representing capacity with a single scalar prevents unnecessary dimensionality growth while maintaining critical constraints.
The aggregate order state is encoded as:
$\boldsymbol{S^{o}_{t}}=(S^{1}_{t},S^{2}_{t},\ldots,S^{2N-1}_{t},S^{2N}_{t}),$
where each aisle $n$ corresponds to two components, $(S^{2n-1}_{t},S^{2n}_{t})$. These values are dynamically computed, capturing the combined item demand weighted by distance, separately for upward and downward traversals within each aisle. As picker location or order demand changes, these reward potentials are continuously updated, ensuring the representation reflects the current system state without explicit spatial detail.
By adopting this feature-based representation, we effectively eliminate redundant fine-grained spatial data, significantly reducing the dimensionality of the state space. The proposed representation retains critical information required for efficient routing decisions—such as aisle-specific potential rewards and capacity constraints—without overwhelming the agent with irrelevant coordinate-level detail. Consequently, this facilitates faster policy learning, improved interpretability, and greater robustness against overfitting to unnecessary spatial complexity.
Appendix B Analysis of the DRL Model with Different Problem Settings
Our DRL model assumes that the picker can initiate movement for order picking from the depot as long as at least one order is pending. However, studies such as Ratliff and Rosenthal (1983) and Lu et al. (2016) establish a higher bound on the initial pick size for picker routing. Furthermore, these studies assume narrow cross-aisles, preventing rerouting through them. This limitation ensures that the picker does not move back in the opposite direction without entering an aisle. In contrast, our approach permits cross-aisle rerouting in the primary experiments, allowing for greater flexibility in picker movement. In this section, we incorporate a combination of these restrictive assumptions into the DRL agent’s environment and evaluate its performance after training under these conditions.
Table 8 summarizes the various DRL models used for comparison. In this section, the parameter $\alpha$ is set to 1. Additionally, we emphasize that the main manuscript includes an analysis of the arrival rate on which the DRL agent should be trained to perform well across a range of arrival rates during testing. However, in this section, we do not repeat that analysis. Instead, we directly apply the insights gained from the main experiments in the manuscript. Specifically, the arrival rate ($\lambda$) used for training each model is chosen based on the best observed performance in terms of average order completion times, as reported in the manuscript’s primary experiments (i.e., the gray-highlighted sections in Table 6). Specifically, for $\lambda\in\{0.01,0.02,0.03,0.04\}$, the DRL model trained on an arrival rate of 0.02 is utilized. For $\lambda=0.05$, the model trained on an arrival rate of 0.04 is selected. Similarly, for $\lambda\in\{0.06,0.07\}$, the model trained on an arrival rate of 0.06 is applied, and for $\lambda\in\{0.08,0.09\}$, the model trained on an arrival rate of 0.08 is used. This approach ensures the selection of models best suited to optimize performance under varying arrival rate conditions, allowing for a focused analysis of the maximum achievable performance across different variations of the DRL model highlighted in Table 8.
Table 8: Variations of the DRL Model Used for Sensitivity Analysis
Model Initial Pick Size Cross-aisle Re-routing
DRL 1 20 No
DRL 2 20 Yes
DRL 3 5 No
DRL 4 5 Yes
DRL 5 1 No
DRL 6 1 Yes
Table 9: Sensitivity Analysis: Average Travel Distance per Order (ATDO)
ATDO (meters) w.r.t. $\lambda$
Model 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09
DRL 1 49.85 52.15 56.58 67.46 54.85 60.89 48.76 69.61 53.54
DRL 2 51.46 54.72 60.03 69.91 58.3 55.82 51.89 45.79 38.88
DRL 3 48.56 53.36 56.32 64.02 59.8 59.36 49.13 69.3 53.41
DRL 4 49.08 54.34 57.16 65.26 61.41 59.35 55.23 46.19 39.59
DRL 5 34.31 41.62 49.56 56.95 60.46 58.74 49.18 68.99 53.01
DRL 6 33.85 40.31 48.7 59.89 63.3 59.03 55.66 46.57 40.0
The results of the experiments are presented in Tables 9, 10, and 11, corresponding to the Average Travel Distance per Order (ATDO), Average Order Completion Time (AOCT), and Percentage of Unfulfilled Orders (PUO) metrics, respectively. These results represent the averages over 10 instances for arrival rates $\lambda=0.01,0.02,\dots,0.09$. As anticipated, the experimental findings suggest that higher initial pick sizes tend to be restrictive, resulting in longer travel distances, increased order completion times, and lower order fulfillment ratios. Consequently, we recommend setting the initial pick size to 1 for this dynamic order-picking scenario. An important question arises: why impose an initial pick size restriction at all? Allowing pickers unrestricted movement, even when no orders are queued, might appear beneficial. However, this could lead to unnecessary empty travels and cause training delays for the DRL agent, particularly under lower order arrival rates. Regarding the performance of the models with initial pick size of unit value, DRL 5 (without cross-aisle rerouting) and DRL 6 (with cross-aisle rerouting) demonstrate comparable results on average. Nonetheless, at higher order arrival rates, the unrestricted variant (DRL 6) generally yields better performance.
Table 10: Sensitivity Analysis: Average Order Completion Time (AOCT)
AOCT (seconds) w.r.t. $\lambda$
Model 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09
DRL 1 480.7 300.8 254.2 254.9 226.3 282.2 295.8 505.0 581.6
DRL 2 474.5 304.5 252.6 254.9 233.4 257.2 299.1 366.6 538.2
DRL 3 333.8 230.1 201.7 213.9 216.8 292.3 307.5 510.0 587.7
DRL 4 329.5 227.6 200.9 214.7 221.8 263.8 321.9 367.5 557.7
DRL 5 66.8 90.1 121.5 164.7 206.5 290.3 305.7 524.2 579.6
DRL 6 67.4 87.2 118.1 171.6 210.2 260.9 323.1 369.0 556.3
Table 11: Sensitivity Analysis: Percentage of Unfulfilled Orders (PUO)
PUO (%) w.r.t. $\lambda$
Model 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09
DRL 1 1.86 1.08 0.95 1.07 0.78 1.02 0.97 1.73 2.16
DRL 2 1.80 1.14 0.84 0.93 0.76 0.68 0.90 1.31 2.00
DRL 3 1.37 0.93 0.98 0.66 0.81 0.94 0.98 1.76 2.08
DRL 4 1.33 0.91 0.92 0.73 0.74 0.74 0.84 1.36 2.19
DRL 5 0.36 0.31 0.43 0.75 0.69 0.86 1.03 1.71 2.01
DRL 6 0.64 0.36 0.30 0.60 0.56 0.69 1.01 1.25 2.27
Appendix C Analysis of the DRL Model with Fluctuating Arrival Rates
In practical warehouse operations, demand rates often fluctuate throughout the operational horizon, characterized by peak and trough periods. Although our proposed DRL framework has been trained and rigorously benchmarked under specific order arrival rates, we also evaluate its performance in scenarios that more accurately reflect real-world dynamics. In this section, we specifically investigate the influence of varying order arrival rates on the proposed DRL model by simulating realistic operational conditions.
To effectively replicate practical variations in demand, we consider two cases with distinct arrival rate sequences: Case 1 ($\{0.02,0.06,0.04,0.08\}$) and Case 2 ($\{0.09,0.01,0.03,0.07\}$). Each arrival rate within a sequence is maintained for two hours, cumulatively representing a full 8-hour work shift. For example, in Case 1, the arrival rate is 0.02 for the first two hours, 0.06 for the next two, followed by 0.04 and 0.08 in subsequent two-hour intervals. During each 2-hour period, we utilize the DRL model previously trained on the arrival rate closest to the observed rate. For instance, when the arrival rate is either 0.01 or 0.02, the DRL model trained at an arrival rate of 0.02 is employed. Tables 12(a) and 12(b) summarize the results of these experiments.
Table 12: Performance metrics for different arrival rate sequences
(a) Arrival rate sequence {0.02, 0.06, 0.04, 0.08}
Model ATDO AOCT PUO
(meters) (seconds) (%)
DRL 60.52 534.80 1.41
Baseline 1 8.19 1,581.31 10.26
Baseline 2 8.30 1,484.38 9.99
Baseline 3 8.29 1,485.79 9.99
Baseline 4 8.29 1,478.38 9.78
Baseline 5 8.29 1,479.00 9.78
(b) Arrival rate sequence {0.09, 0.01, 0.03, 0.07}
Model ATDO AOCT PUO
(meters) (seconds) (%)
DRL 68.50 415.10 0.91
Baseline 1 8.21 565.71 1.42
Baseline 2 8.82 467.51 1.22
Baseline 3 8.82 462.46 1.22
Baseline 4 8.86 466.82 1.24
Baseline 5 8.86 464.02 1.21
The experimental results demonstrate the adaptability of the DRL models under dynamically changing order arrival conditions, highlighting their robustness compared to traditional baseline models. In Case 1, where the arrival rate sequence follows {0.02, 0.06, 0.04, 0.08}, the baseline models perform poorly, with significantly higher throughput times and a much larger percentage of unfulfilled orders compared to the DRL model. The DRL model achieves a 66.2% reduction in throughput time and a 85.6% decrease in unfulfilled orders relative to the best-performing baseline. In Case 2, where the arrival rate sequence is {0.09, 0.01, 0.03, 0.07}, the baseline models exhibit relatively improved performance. This can be attributed to the highest arrival rate occurring in the initial phase, leading to a backlog of unfulfilled orders that can be processed during the subsequent phases when order arrivals decrease. However, even in this scenario, the DRL model continues to outperform the baselines, achieving a 10.2% reduction in throughput time and a 24.8% decrease in unfulfilled orders compared to the best-performing baseline.
These findings underscore the effectiveness of the proposed DRL approach in handling fluctuating demand patterns, ensuring lower order fulfillment delays and improved operational efficiency across different demand scenarios.